OpenStudy (anonymous):

how to find the equation of diameter..? of a circle..whose equation is x2+Y2-12x+4Y+6=0

razor99 (razor99):

could u explain it clearly

OpenStudy (anonymous):

You need to complete the square for both the x and y values within the equation.

OpenStudy (anonymous):

mark..we have to find the equation of diameter?

OpenStudy (anonymous):

can anyof u tell me how to find the equation of diameter?

OpenStudy (anonymous):

You need to put the equation in standard form first. \[x^2+y^2-12x+4y+6=0\] \[x^2-12x+y^2+4y+6=0\] \[(x^2-12x)+(y^2+4y)+6=0\] \[(x^2-12x +36-36)^2+(y^2+4y+4-4)+6=0\] this is the completing the square part \[(x^2-12x +36)^2-36+(y^2+4y+4)-4+6=0\] \[(x^2-12x +36)^2+(y^2+4y+4)+38=0\] \[(x-6)^2+(y+2)^2+38=0\]

OpenStudy (anonymous):

oops, thte +38 should be -34

OpenStudy (anonymous):

Once you move the -34 over, it means the radius is sqrt(34)

razor99 (razor99):

okay here is the formulae for the diameter

razor99 (razor99):

hey is it area or perimeter

OpenStudy (anonymous):

its diameter of circle

OpenStudy (anonymous):

yeah so the diameter is 2x the square root of 34

razor99 (razor99):

okay formulae of a diameter d=2xr

OpenStudy (anonymous):

but we hav to find the equation not only diameter

razor99 (razor99):

oh first we have to slove it right

OpenStudy (anonymous):

hmm..yup

OpenStudy (anonymous):

hey razor r u on fb?

razor99 (razor99):

yeah

OpenStudy (anonymous):

wud u lyk 2 b mah frnd..?

razor99 (razor99):

okay but i am a boy

OpenStudy (anonymous):

uhh..y r u tellng mee lolz..!! it dsnt seem scary 2 mee

razor99 (razor99):

oh lets talk some other time okay u have a equation to do

OpenStudy (anonymous):

yea..i got it

razor99 (razor99):

dont worry jimmyrep is coming

OpenStudy (anonymous):

whoz he?

razor99 (razor99):

comeon u would know

razor99 (razor99):

Equation x2+Y2-12*x+4*Y+6=0 solved for x x = (Y2+4*Y+x2+6)/12

razor99 (razor99):

so the answer is

razor99 (razor99):

Equation x2+Y2-12*x+4*Y+6=0 solved for x x = (Y2+4*Y+x2+6)/12